# Missing Pieces: How Do Designs that Expose Uncertainty Longitudinally Impact Trust in AI Decision Aids? An In Situ Study of Gig Drivers

### v2

@@ B.5. Results
P1, P2, and P3 all felt that it was better for the tool to have a simple, easy-to-use design. All three indicated that the prototype fulfilled this requirement, although P3 suggested that the wording and design improvements would be necessary (in particular, they felt that the monotone colour scheme of the tool was confusing). P2 and P3 felt that the design needed to be mobile-friendly. The participants also mentioned other desiderata:
@@ C.3. Feedback on Schedules
  • How did your response to the recommended schedules change throughout the week, if at all?
@@ Appendix D Full Quantitative Results
D.1. Daily Trust and Reliance Statistics
@@
In Table 3, we report the means, standard errors, and 95% confidence intervals of the daily trust and reliance scores plotted in Figures 4 and 6. We show both the overall statistics as well as the statistics for each condition. In Figure 6, we show visualisations for daily reliance scores omitted from the main text.
D.2. Trust and Reliance Score LMMs
@@ D.3. Post-Interaction Trust
We fitted an ordinary least squares (OLS) models for the post_trust_score using statsmodels. This model included the condition along with all previous trust (pre_trust_score and daily trust_score) and reliance measurements.
@@
For Conditions (B)/(D)/(R)/(RH), the mean post-interaction trust scores were 4.000, 3.720, 3.857, and 3.618. None of these conditions were significantly different from each other, and the coefficients for previous trust and reliance scores were not statistically significant either. The questions we adapted from the TiA asked participants to consider the entire duration of their interaction with the schedule recommendation tool. As a result, this broad, retrospective reflection may have failed to capture more nuanced longitudinal changes in trust and reliance like those we described in Section 5.
@@ Appendix E Figma Prototype Design
Figure 7. Screenshot of Figma prototype for the constraint page. On behalf of participants, the interviewer clicked on radio buttons, input fields, and table cells to set constraints. For technical reasons, input fields for the numerical constraints were implemented as dropdown menus.

### v4

@@ B.5. Results
P1, P2, and P3 all felt that it was better for the tool to have a simple, easy-to-use design. All three indicated that the prototype fulfilled this requirement, although P3 suggested that wording and design improvements would be necessary (in particular, they felt that the monotone colour scheme of the tool was confusing). P2 and P3 felt that the design needed to be mobile-friendly. The participants also mentioned other desiderata:
@@ C.3. Feedback on Schedules
  • How did your response to recommended schedules change throughout the week, if at all?
@@ Appendix D Full Quantitative Results
D.1. Trust and Reliance Score LMMs
D.2. Daily Trust and Reliance Statistics
@@
[…]
On the left, a line plot with seven points decreases from 0.176 on Day 1 to 0.088 on Day 6. The general trend is downwards but not monotonic. 95% confidence intervals for each of the points are overlapping. On the right, a boxplot shows two boxes for each day: one red (perceived inaccurate) and one green (perceived accurate). There is no clear trend in the IQRs, which all overlap. Most of the IQRs are between 0 and 1, but some of them consist of 0 only.
In Figure 6, we show visualisations for daily reliance scores omitted from the main text. In Table 3, we report the means, standard errors, and 95% confidence intervals of the daily trust and reliance scores plotted in Figures 4 and 6. We show both the overall statistics as well as the statistics for each condition.
@@ D.3. Post-Interaction Trust
We fitted an ordinary least squares (OLS) model for the post_trust_score using statsmodels. This model included the condition along with all previous trust (pre_trust_score and daily trust_score) and reliance measurements.
@@
For Conditions (B)/(D)/(R)/(RH), the mean post-interaction trust scores were 4.000, 3.720, 3.857, and 3.618. None of these conditions were significantly different from each other, and the coefficients for previous trust and reliance scores were not statistically significant either. The questions we adapted from the TiA asked participants to consider the entire duration of their interaction with the schedule recommendation tool. This broad, retrospective reflection may have failed to capture more nuanced longitudinal changes in trust and reliance like those we described in Section 5.
@@ Appendix E Figma Prototype Design
Figure 7. Screenshot of Figma prototype for the constraint page. On behalf of participants, the interviewer clicked on radio buttons, input fields, and table cells to set their constraints. For technical reasons, input fields for the numerical constraints were implemented as dropdown menus.
[…]
Screenshot of a webpage titled ”Shift Recommendation Tool”. The constraint page is shown, with various inputs being set to mock values. In response to ”Would you like to set a maximum number of hours that you would like to drive to earn as much money as possible, or a minimum amount of money that you would like to earn in as few hours as possible?”, the latter is selected. In response to ”Would you like to set these constraints for each day individually, or for the entire week?”, the former is selected. In response to ”What is the minimum amount of money you would like to earn for each day?”, \$140-\$150 is entered for each day of the week. In response to ”When are you available during the week?”, time slots corresponding to 7 am to 2 pm on every day of the week are selected (as indicated by dark cells in a tabular schedule).
@@
[…]
Screenshot of a webpage titled ”Shift Recommendation Tool”. The schedule page is shown with mock estimates. The top of the schedule page reads, ”Based on historical data, it is estimated that you will earn a weekly average of”. This is followed by a table that reads, ”On an average week: \$593”, ”On a bad week: \$152”, and ”On a good week: \$1135”. Further below, a tabular schedule is shown for each day of the week. Each cell reads, ”Average”, followed by the estimated hourly earning in bold. Recommended time slots are indicated by dark cells, and non-recommended time slots are indicated by light cells.
@@ Appendix F Final Tool Design
[…]
Screenshot of a webpage titled ”Tool Interaction”. Below a set of instructions, the constraint page is shown in a grey iframe. Various inputs for constraints are set to values entered by interview participant P1. In response to ”Would you like to specify a maximum number of hours that you would like to drive, or would you prefer to specify a minimum amount of money that you would like to earn?”, P1 selected the latter. In response to ”Would you like to set these constraints for each day individually, or for the entire week?”, P1 selected the latter. In response to ”What is the minimum amount of money you would like to earn for the entire week?”, P1 entered \$1400. In response to ”When are you available during the week?”, P1 selected time slots (as indicated by blue cells in a tabular schedule) corresponding to 3 am to 7 am, 8 am to 4 pm, and 10 pm to 12 am on every day of the week.
@@ F.1. Schedule Page Conditions
[…]
Screenshot of a webpage titled ”Tool Interaction”. Below a set of instructions, the schedule page in Condition (B) for interview participant P1 is shown in a grey iframe. The top of the schedule page reads, ”Based on historical data, it is estimated that you will earn a weekly average of”, followed by an estimate of \$1412 below. Further below, a tabular schedule is shown for each day of the week. Each cell reads, ”Average”, followed by the estimated hourly earning in bold. Recommended time slots are indicated by blue cells, non-recommended time slots during which P1 is still available are indicated by dark text, and time slots during which P1 is not available are indicated by light text.
@@
[…]
Screenshot of a webpage titled ”Tool Interaction”. Below a set of instructions, the schedule page in Condition (D) for interview participant P1 is shown in a grey iframe. The top of the schedule page reads, ”Based on historical data, it is estimated that you will earn”. Unlike Condition (B), this is followed by a table that reads, ”An average of”, alongside estimates for each day: \$210 for Monday, \$194 for Tuesday, \$175 for Wednesday, \$173 for Thursday, \$194 for Friday, \$226 for Saturday, and \$237 for Sunday. Further below, a tabular schedule is shown for each day of the week. Each cell reads, ”Average”, followed by the estimated hourly earning in bold. Recommended time slots are indicated by blue cells, non-recommended time slots during which P1 is still available are indicated by dark text, and time slots during which P1 is not available are indicated by light text.
@@
[…]
Screenshot of a webpage titled ”Tool Interaction”. Below a set of instructions, the schedule page in Condition (R) for interview participant P1 is shown in a grey iframe. The top of the schedule page reads, ”Based on historical data, it is estimated that you will earn”. Unlike Conditions (B) and (D), this is followed by a table that reads, ”On an average week: \$1412”, ”On a bad week: \$391”, and ”On a good week: \$3032”. Further below, a tabular schedule is shown for each day of the week. Each cell reads, ”Average”, followed by the estimated hourly earning in bold. Unlike Conditions (B) and (D), this is also followed by ”Low”, with the 10th percentile of earnings, and ”High”, with the 90th percentile of earnings. Recommended time slots are indicated by blue cells, non-recommended time slots during which P1 is still available are indicated by dark text, and time slots during which P1 is not available are indicated by light text.
@@
[…]
Screenshot of a webpage titled ”Tool Interaction”. Below a set of instructions, the schedule page in Condition (RH) for interview participant P1 is shown in a grey iframe. Unlike Conditions (B), (D), and (R), the top of the schedule page reads, ”In general, every driver may encounter different conditions while they are driving, which leads to differences in earnings from individual to individual and from one day to the next. On average, based on historical data, a driver following this schedule will earn”. Also unlike Conditions (B) and (D), this is followed by a table that reads, ”On an average week: \$1412”, ”On a bad week: \$391”, and ”On a good week: \$3032”. Further below, a tabular schedule is shown for each day of the week. Each cell reads, ”Average”, followed by the estimated hourly earning in bold. Unlike Conditions (B) and (D), this is also followed by ”Low”, with the 10th percentile of earnings, and ”High”, with the 90th percentile of earnings. Recommended time slots are indicated by blue cells, non-recommended time slots during which P1 is still available are indicated by dark text, and time slots during which P1 is not available are indicated by light text.